OpenStudy (anonymous):

Solve by factoring: x2-8x+15=0

OpenStudy (anonymous):

x^2 - 8x + 15 = 0 (factor by finding two integers that multiply to equal 15, but add to equal -8) (x - 5)(x - 3) = 0 (zero-product property) x - 5 = 0 or x - 3 = 0 x = 5 or x = 3 <===ANSWER

OpenStudy (anonymous):